# Reducing leakage of single-qubit gates for superconducting quantum processors using analytical control pulse envelopes

### v1

@@ Abstract
Improving the speed and fidelity of quantum logic gates is essential to reach quantum advantage with future quantum computers. However, fast logic gates lead to increased leakage errors in superconducting quantum processors based on qubits with low anharmonicity, such as transmons. To reduce leakage errors, we propose and experimentally demonstrate two new analytical methods, Fourier ansatz spectrum tuning derivative removal by adiabatic gate (FAST DRAG) and higher-derivative (HD) DRAG, both of which enable shaping single-qubit control pulses in the frequency domain to achieve stronger suppression of leakage transitions compared to previously demonstrated pulse shapes. Using the new methods to suppress the $ef$-transition of a transmon qubit with an anharmonicity of -212 MHz, we implement $R_{X}(\pi/2)$-gates with a leakage error below $3.0\times 10^{-5}$ down to a gate duration of 6.25 ns, which corresponds to a 20-fold reduction in leakage compared to a conventional Cosine DRAG pulse. Employing the FAST DRAG method, we further achieve an error per gate of $(1.56\pm 0.07)\times 10^{-4}$ at a 7.9-ns gate duration, outperforming conventional pulse shapes both in terms of error and gate speed. Furthermore, we study error-amplifying measurements for the characterization of temporal microwave control pulse distortions, and demonstrate that non-Markovian coherent errors caused by such distortions may be a significant source of error for sub-10-ns single-qubit gates unless corrected using predistortion.
@@ I Introduction
Superconducting qubits provide a promising platform for realizing large-scale quantum computers as they enable high-fidelity quantum logic gate operations in the nanosecond regime [1, 2, 3, 4, 5, 6], potentially resulting in significantly faster computations compared to other state-of-the-art quantum computing platforms, such as neutral atoms [7, 8, 9] or trapped ions [10, 11]. Improving the gate speed further would not only increase the clock speed of the quantum computer but it would also reduce the incoherent gate errors caused by decoherence, thus enabling even higher fidelities. Currently, the large-scale superconducting quantum processing units (QPU) are based on transmon qubits which suffer from a low anharmonicity typically in the range of $|\alpha/(2\pi)|\in[150,300]$ MHz [1, 2, 12]. Thus, fast logic gates cause the frequency spectra of the control pulses to overlap with transitions out of the computational subspace – such as the $ef$-transition – resulting in leakage errors [13, 14, 15, 16]. Leakage is particularly harmful for quantum error correction applications [17, 18, 19] in transmon-based processors due to the generation of non-local errors caused by the entanglement between the leakage states [20].
@@
Several approaches have been proposed and demonstrated to mitigate leakage errors of quantum logic gates implemented with superconducting qubits [13, 14, 15, 16, 21, 22, 23, 24, 25, 26, 27, 1]. Leakage errors caused by single-qubit gate control pulses are often reduced using the derivative removal by adiabatic gate (DRAG) method that sets the quadrature envelope of a control pulse equal to a scaled derivative of its in-phase envelope [13, 15, 16]. This effectively suppresses the frequency spectrum of the control pulse at a specific frequency, such as the $ef$-transition frequency, reducing leakage related to the suppressed transition. Additionally, there exist leakage reduction methods based on shaping the control pulse envelopes, an example of which is the Slepian pulse shape [24, 5, 4] used especially for two-qubit gates to minimize the spectral energy above a given cutoff frequency. Furthermore, optimal control methods have been shown to reduce leakage errors of fast gates. For example, closed-loop optimization of piecewise constant 4.16-ns $R_{X}(\pi/2)$ control pulses with 23 parameters resulted in a leakage error per Clifford gate of $4.4\times 10^{-4}$ on a transmon when using the experimental gate fidelity as the cost function [26]. However, such closed-loop optimization methods may require impractically long calibration times of up to tens of hours.
@@
In this work, we demonstrate two new methods for engineering single-qubit gate pulse shapes that enable a more flexible control of the frequency spectrum resulting in significantly reduced leakage errors for fast gates without the need for closed-loop optimization. Applying the new methods to implement single-qubit gates on a transmon with an anharmonicity of $\alpha/(2\pi)=-212$ MHz, we experimentally achieve an average leakage error per $R_{X}(\pi/2)$-gate below $3\times 10^{-5}$ down to a gate duration of $t_{\mathrm{g}}=6.25$ ns corresponding to a 20-fold reduction in leakage compared to DRAG with a conventional cosine envelope. Importantly, one of the novel approaches presented here, Fourier ansatz spectrum tuning (FAST) DRAG, enables the implementation of a 7.9-ns $R_{X}(\pi/2)$-gate with a gate error of $(1.56\pm 0.07)\times 10^{-4}$, reducing the gate error and improving the gate speed over the conventional pulse shapes. Furthermore, we study error mechanisms relevant for fast single-qubit gates with fidelities approaching 99.99%, and find that non-Markovian coherent errors caused by microwave control pulse distortions [28] may be a major error source in addition to incoherent errors and leakage. As the measurement of the full process tensor [29, 30, 31] capturing non-Markovian errors is both time consuming and often associated with difficulties in connecting the results to a realistic error model, we instead study simple error-amplification experiments that characterize coherent errors caused by such pulse distortions and enable the calibration of parameters needed for control pulse predistortion.
@@ II.1 Introduction to single-qubit gates and DRAG
Single-qubit rotations of superconducting qubits are typically implemented with nearly resonant microwave drive pulses as illustrated in Fig. 1(a) with the drive pulses applied to a drive line which is either capacitively or inductively coupled to the qubit. In the frame rotating at the microwave drive frequency $f_{\mathrm{d}}=\omega_{\mathrm{d}}/(2\pi)$, the Hamiltonian of a driven weakly non-linear oscillator, such as a transmon qubit, can be written as [13]
@@
Conventionally, single-qubit control pulses are shaped with DRAG to reduce either phase errors – which we here refer to as DRAG-P – or both phase and leakage errors, which we refer to as DRAG-L [15, 16]. For both of the DRAG variants, the quadrature envelope is given by $\Omega_{\mathrm{Q}}(t)=-\beta\dot{\Omega}_{I}(t)/\alpha$, where $\beta$ denotes the DRAG coefficient, and $\Omega_{\mathrm{I}}(t)$ is the in-phase pulse envelope, which can be, e.g., a Gaussian with $\Omega_{\mathrm{I}}(t)=\mathrm{exp}[-(t-t_{\mathrm{p}}/2)^{2}/(2\sigma^{2})]-\mathrm{exp}[-t_{\mathrm{p}}^{2}/(8\sigma^{2})]$ or a raised cosine with $\Omega_{\mathrm{I}}(t)=[1-\cos(2\pi t/t_{\mathrm{p}})]/2$ as shown in Fig. 1(b). In the DRAG-P variant, the DRAG coefficient $\beta$ is tuned to cancel the AC Stark shift during the gate, which for a transmon occurs at $\beta\approx 0.5$ [32, 22]. For sufficiently slow gates, the leakage errors can be ignored.
In the second variant, DRAG-L, the DRAG coefficient $\beta$ is tuned such that the spectrum of the control pulse has a zero crossing in the vicinity of the $ef$-transition frequency, which occurs at $\beta\approx 1.0$. This spectral suppression leads to a lower leakage error, thus potentially enabling faster and more accurate gates compared to DRAG-P. However, the choice of $\beta\approx 1.0$ is not optimal for eliminating the AC Stark shift, and thus phase errors need to be corrected with another approach, such as by detuning the drive frequency during the gate [15] or by applying virtual-$Z$ rotations, i.e., updating the phase of the microwave drive, before and after each gate [16]. If the phase errors are accounted for by tuning the drive frequency during the gate, the optimal values of the drive detuning and $\beta$ are coupled, which complicates calibration. Here, we use virtual-$Z$ rotations that are especially convenient for gate sets based on $R_{X}(\pi/2)$ single-qubit rotations and two-qubit $CZ$ rotations. With this gate set, the implementation of a $R_{X}(\pi)$-gate, however, requires two separate $R_{X}(\pi/2)$-gates because the qubit state cannot be evolved from pole to pole without applying virtual Z-rotations at some point between the poles. Despite of lower errors provided by DRAG-L, the DRAG-P variant has been more commonly used in literature due to easier calibration and the dominance of phase errors over leakage in early superconducting qubits [32, 21, 35, 36, 37].
@@ II.2 Control pulses based on Fourier ansatz spectrum tuning
For short microwave pulses or in the presence of multiple leakage transitions, it would be desirable to suppress the frequency spectrum across one or many frequency intervals of adjustable width, which is not possible with conventional DRAG pulses suppressing the spectrum across a single narrow frequency range. Importantly, the frequencies of the leakage transitions have finite widths induced by dephasing [34] and they may shift during short gates due to the above-mentioned AC Stark shifts. Thus, it may be beneficial to suppress the spectrum of the control pulse across wider frequency intervals around leakage transitions to reduce the leakage errors of fast gates beyond the performance offered by conventional DRAG pulses.
To overcome these issues, we present a novel pulse shaping method that parametrizes a control pulse envelope in terms of one or multiple frequency intervals, across which the spectral energy is minimized to reduce leakage. We call the method Fourier ansatz spectrum tuning (FAST) since the spectral shaping is achieved using a control pulse envelope expressed as a Fourier cosine series
@@
where $g_{n}(t)=1-\cos(2\pi nt/t_{\mathrm{p}})$ is the $n$th basis function parametrized to ensure continuity of the cosine series similar to Ref. [25]. Furthermore, $t_{\mathrm{p}}$ denotes the pulse duration, $A$ is the overall amplitude, $N$ is the number of cosine terms, and $\{c_{n}\}$ are the set of Fourier coefficients.
@@
where $\bm{A}_{nm}=\sum_{j=1}^{k}w_{j}\int_{f_{\mathrm{l},j}}^{f_{\mathrm{h},j}}\hat{g}_{n}(f)\hat{g}_{m}^{*}(f)\mathrm{d}f$, $\bm{b}=(1,\dots,1)^{T}\in\mathbb{R}^{N\times 1}$, and $\mu$ is the Lagrangian multiplier. A more detailed derivation is provided in Appendix A. The FAST method can be viewed as an extension of the Slepian [24], and it parametrizes the control pulse in the frequency domain, while providing a mapping to the time domain parameters $\{c_{n}\}$ using Eq. (5).
@@
To implement fast, low-leakage single-qubit gates, we combine the FAST method with DRAG by using FAST to shape the in-phase envelope of the control pulse and applying DRAG to obtain the quadrature component. The resulting FAST DRAG pulse illustrated in Fig. 1(b) has multiple benefits compared to using DRAG with traditional pulse envelopes, such as the raised cosine. First, the FAST DRAG pulse enables a stronger and wider spectral suppression of the $ef$-transition compared to conventional approaches as shown in Fig. 1(d), which helps to mitigate leakage to the $f$-state. To achieve this, we apply the FAST method to suppress the spectrum of the in-phase component across a frequency interval $[f_{\mathrm{l},1},f_{\mathrm{h},1}]\ni|\alpha|/(2\pi)$, and use DRAG-L with $\beta\approx 1$ to further suppress the Fourier transform $\hat{\Omega}_{\mathrm{IQ}}$ of the complex envelope that is of the product form $\hat{\Omega}_{\mathrm{IQ}}(f)=[1-2\pi\beta f/\alpha]\times\hat{\Omega}_{I}(f)$. The strong suppression of $\hat{\Omega}_{\mathrm{IQ}}(f)$ around $\alpha/(2\pi)$ then leads to a low spectral amplitude $|\hat{\Omega}_{\mathrm{RF}}(f)|$ around the $ef$-transition frequency assuming $\omega_{\mathrm{d}}=\omega_{\mathrm{q}}$. In addition to suppressing the signal around the $ef$-transition, we use a second frequency interval $[f_{\mathrm{l},2},f_{\mathrm{h},2}]$ to reduce the spectral energy above a given cutoff frequency $f_{\mathrm{c}}=f_{\mathrm{l},2}$. This favorably reduces the bandwidth of the control pulse and often indirectly lowers the peak amplitude, thus effectively regularizing the pulse shape. As a further benefit, it is relatively straightforward to calibrate FAST DRAG pulses since the parameters $A$, $\beta$, and $\{[f_{\mathrm{l},j},f_{\mathrm{h},j}]\}$ are mostly decoupled, and the the locations of the suppressed frequency intervals $\{[f_{\mathrm{l},j},f_{\mathrm{h},j}]\}$ do not affect the effective drive frequency due to a symmetric suppression of the spectrum (see Fig. 1(c)).

### v2

@@ Abstract
Improving the speed and fidelity of quantum logic gates is essential to reach quantum advantage with future quantum computers. However, fast logic gates lead to increased leakage errors in superconducting quantum processors based on qubits with low anharmonicity, such as transmons. To reduce leakage errors, we propose and experimentally demonstrate two new analytical methods, Fourier ansatz spectrum tuning derivative removal by adiabatic gate (FAST DRAG) and higher-derivative (HD) DRAG, both of which enable shaping single-qubit control pulses in the frequency domain to achieve stronger suppression of leakage transitions compared to previously demonstrated pulse shapes. Using the new methods to suppress the $ef$-transition of a transmon qubit with an anharmonicity of -212 MHz, we implement $R_{X}(\pi/2)$-gates achieving a leakage error below $3.0\times 10^{-5}$ down to a gate duration of 6.25 ns without the need for iterative closed-loop optimization. The obtained leakage error represents a 20-fold reduction in leakage compared to a conventional Cosine DRAG pulse. Employing the FAST DRAG method, we further achieve an error per gate of $(1.56\pm 0.07)\times 10^{-4}$ at a 7.9-ns gate duration, outperforming conventional pulse shapes both in terms of error and gate speed. Furthermore, we study error-amplifying measurements for the characterization of temporal microwave control pulse distortions, and demonstrate that non-Markovian coherent errors caused by such distortions may be a significant source of error for sub-10-ns single-qubit gates unless corrected using predistortion.
@@ I Introduction
Superconducting qubits provide a promising platform for realizing large-scale quantum computers as they enable high-fidelity quantum logic gate operations in the nanosecond regime [1, 2, 3, 4, 5, 6, 7, 8], potentially resulting in significantly faster computations compared to other state-of-the-art quantum computing platforms, such as neutral atoms [9, 10, 11] or trapped ions [12, 13]. Improving the gate speed further would not only increase the clock speed of the quantum computer but it would also reduce the incoherent gate errors caused by decoherence, thus enabling even higher fidelities. Currently, the large-scale superconducting quantum processing units (QPU) are based on transmon qubits which suffer from frequency crowding and a low anharmonicity typically in the range of $|\alpha/(2\pi)|\in[150,300]$ MHz [1, 2, 14]. Thus, fast logic gates cause the frequency spectra of the control pulses to overlap with transitions out of the computational subspace – such as the $ef$-transition – resulting in leakage errors [15, 16, 17, 18]. Leakage is particularly harmful for quantum error correction applications [19, 20, 21] in transmon-based processors due to the generation of non-local errors as a result of leakage being propagated by entangling two-qubit gates [22].
@@
There are two main strategies to mitigate leakage errors in superconducting qubits. The first strategy aims to prevent the occurrence of leakage in the first place, e.g., using pulse shape engineering [15, 16, 17, 18, 23, 24, 25, 26, 27, 28, 7, 1, 29, 30, 31], whereas the second strategy relies on leakage reduction units converting already occurred leakage errors into Pauli errors [22, 32, 33]. In this work, we focus on the first strategy that enables the reduction of the overall error for fast logic operations. Leakage errors caused by single-qubit control pulses are often suppressed using the derivative removal by adiabatic gate (DRAG) method that sets the quadrature envelope of a control pulse equal to a scaled derivative of its in-phase envelope [15, 17, 18]. This effectively suppresses the frequency spectrum of the control pulse at a specific frequency, such as the $ef$-transition frequency, reducing leakage related to the suppressed transition. Additionally, the leakage error can be reduced by the choice of the control pulse envelopes, for example, using the Slepian pulse shape [26, 5, 4] that is commonly applied in two-qubit gates to minimize the spectral energy above a given cutoff frequency. Furthermore, optimal control methods have been shown to reduce leakage errors of fast gates. For example, the lowest leakage error for fast transmon-based single-qubit gates has been achieved using experimental closed-loop optimization of piecewise constant $R_{X}(\pi/2)$ control pulses with 23 parameters, which resulted in a leakage rate per Clifford gate of $4.4\times 10^{-4}$ at a pulse duration of $t_{\mathrm{g}}=4.16$ ns ($t_{\mathrm{g}}|\alpha|/(2\pi)\approx 1.31$) [7]. However, such closed-loop optimization methods may require a large number of iterations, leading to increased calibration time [7]. Open-loop optimal control techniques may provide significantly faster convergence by optimizing the control pulses in a simulation [34, 35, 36, 37, 38, 39], but the obtained pulse shapes may not transfer well to an experimental implementation if the assumed model Hamiltonian does not capture all the relevant details of the experimental system, such as electronics limitations, pulse distortions or parasitic interactions.
@@
Two-qubit gate fidelities in superconducting quantum processors have recently improved considerably [5, 40, 4, 6, 41, 42], and hence, further improvements are needed for single-qubit gates that typically appear in quantum circuits in higher numbers compared to two-qubit gates [43, 40, 14]. In this work, we demonstrate two new methods for engineering single-qubit gate pulse shapes that enable a more flexible control of the frequency spectrum resulting in significantly reduced leakage errors for fast gates without the need for closed-loop optimization. Applying the new methods to implement single-qubit gates on a transmon with an anharmonicity of $\alpha/(2\pi)=-212$ MHz, we experimentally achieve an average leakage error per $R_{X}(\pi/2)$-gate below $3\times 10^{-5}$ down to a gate duration of $t_{\mathrm{g}}=6.25$ ns ($t_{\mathrm{g}}|\alpha|/(2\pi)\approx 1.33$) approaching the speed limit of single-qubit gates $t_{\mathrm{g}}\sim 2\pi/|\alpha|$ [27, 44]. At such short gate durations, the proposed methods reduce the leakage error by a factor of 20 compared to DRAG with a conventional cosine envelope and by an order of magnitude compared to the previous lowest leakage rate at a comparable gate duration [7]. Importantly, one of the novel approaches presented here, Fourier ansatz spectrum tuning (FAST) DRAG, enables the implementation of a 7.9-ns $R_{X}(\pi/2)$-gate with a gate error of $(1.56\pm 0.07)\times 10^{-4}$, reducing the gate error and improving the gate speed over the conventional pulse shapes. In addition to low errors, the proposed methods also reduce the required peak power of sub-10-ns single-qubit gates compared to conventional pulse shapes. Furthermore, we study error mechanisms relevant for fast single-qubit gates with fidelities approaching 99.99%, and find that non-Markovian coherent errors caused by microwave control pulse distortions [45] may be a major error source in addition to incoherent errors and leakage. We characterize the microwave pulse distortions in our experimental system using simple error amplification experiments and show that the gate fidelity can be improved by predistorting the microwave control pulses.
@@ II.1 Introduction to single-qubit gates and DRAG
Single-qubit rotations of superconducting qubits are typically implemented with nearly resonant microwave drive pulses as illustrated in Fig. 1(a) with the drive pulses applied to a drive line which is either capacitively or inductively coupled to the qubit. In the frame rotating at the microwave drive frequency $f_{\mathrm{d}}=\omega_{\mathrm{d}}/(2\pi)$, the Hamiltonian of a driven non-linear oscillator, such as a transmon qubit, can be written as [15]
@@
Conventionally, single-qubit control pulses are shaped with DRAG to reduce either phase errors – which we here refer to as DRAG-P – or both phase and leakage errors, which we refer to as DRAG-L [17, 18]. For both of the DRAG variants, the quadrature envelope is given by $\Omega_{\mathrm{Q}}(t)=-\beta\dot{\Omega}_{I}(t)/\alpha$, where $\beta$ denotes the DRAG coefficient, and $\Omega_{\mathrm{I}}(t)$ is the in-phase pulse envelope, which can be, e.g., a Gaussian with $\Omega_{\mathrm{I}}(t)=\mathrm{exp}[-(t-t_{\mathrm{p}}/2)^{2}/(2\sigma^{2})]-\mathrm{exp}[-t_{\mathrm{p}}^{2}/(8\sigma^{2})]$ or a raised cosine with $\Omega_{\mathrm{I}}(t)=[1-\cos(2\pi t/t_{\mathrm{p}})]/2$, as shown in Fig. 1(b). In the DRAG-P variant, the DRAG coefficient $\beta$ is tuned to cancel the AC Stark shift during the gate, which for a transmon occurs at $\beta\approx 0.5$ [46, 24]. For sufficiently slow gates, the leakage errors can be ignored.
In the second variant, DRAG-L, the DRAG coefficient $\beta$ is tuned such that the spectrum of the control pulse has a zero crossing in the vicinity of the $ef$-transition frequency, which occurs at $\beta\approx 1.0$. This spectral suppression leads to a lower leakage error, thus potentially enabling faster and more accurate gates compared to DRAG-P. Despite lower errors provided by DRAG-L, the DRAG-P variant has been more commonly used in literature due to easier calibration and the dominance of phase errors over leakage in early superconducting qubits [46, 23, 49, 50, 51]. Namely, the choice of $\beta\approx 1.0$ is not optimal for eliminating the AC Stark shift, and thus phase errors need to be corrected with another approach for the DRAG-L variant, such as by detuning the drive frequency during the gate [17] or by applying virtual Z-rotations, i.e., updating the phase of the microwave drive, before and after each gate [18]. If the phase errors are accounted for by tuning the drive frequency during the gate, the optimal values of the drive detuning and $\beta$ are coupled, which complicates calibration. On the other hand, virtual Z-rotations are straightforward to use with gate sets based on $R_{X}(\pi/2)$-rotations and two-qubit gates commuting with $\hat{\sigma}_{z}$. However, the implementation of a $R_{X}(\pi)$-gate requires two separate $R_{X}(\pi/2)$-gates as the qubit state cannot be evolved from pole to pole without applying virtual Z-rotations at some point between the poles. In principle, arbitrary-angle single-qubit gates can be decomposed with at most two $R_{X}(\pi/2)$-gates and three virtual Z-rotations [18]. In this work, we decompose the single-qubit Clifford group using $R_{X}(\pi/2)$-gates as explained in Sec. III and use virtual Z-rotations to mitigate phase errors of DRAG-L pulses.
@@ II.2 Control pulses based on Fourier ansatz spectrum tuning
For short microwave pulses or in the presence of multiple leakage transitions, it would be desirable to suppress the frequency spectrum across one or many frequency intervals of adjustable width, which is not possible with conventional DRAG pulses suppressing the spectrum across a single narrow frequency range. Importantly, the frequencies of the leakage transitions may shift during short gates due to AC Stark shifts or they may have finite widths induced by dephasing [48]. Thus, it may be beneficial to suppress the spectrum of the control pulse across wider frequency intervals around leakage transitions to reduce the leakage errors of fast gates beyond the performance offered by conventional DRAG pulses.
To overcome these issues, we present a novel pulse shaping method that parametrizes a control pulse envelope in terms of one or multiple configurable frequency intervals, across which the spectral energy is minimized to reduce leakage. We call the method Fourier ansatz spectrum tuning (FAST) since the spectral shaping is achieved using a control pulse envelope expressed as a Fourier cosine series
@@
where $g_{n}(t)=[1-\cos(2\pi nt/t_{\mathrm{p}})]\Pi(t/t_{\mathrm{p}}-1/2)$ is the $n$th basis function parametrized to ensure continuity of the cosine series similar to Ref. [27]. Furthermore, $A$ is the overall amplitude, $N$ is the number of cosine terms, $\{c_{n}\}$ are the set of Fourier coefficients, $t_{\mathrm{p}}$ denotes the pulse duration, and $\Pi(x)$ is the rectangular window function between $x\in[-1/2,1/2]$, which sets the pulse envelope to 0 outside of $[0,t_{\mathrm{p}}]$.
@@
where $\bm{A}_{nm}=\sum_{j=1}^{k}w_{j}\int_{f_{\mathrm{l},j}}^{f_{\mathrm{h},j}}\hat{g}_{n}(f)\hat{g}_{m}^{*}(f)\mathrm{d}f$, $\bm{b}=(1,\dots,1)^{T}\in\mathbb{R}^{N\times 1}$, $\bm{0}=(0,\dots,0)^{T}\in\mathbb{R}^{N\times 1}$, and $\mu$ is the Lagrangian multiplier. A more detailed derivation is provided in Appendix A. The FAST method can be viewed as an extension of the Slepian [26], and it parametrizes the control pulse in the frequency domain, while providing a mapping to the time-domain parameters $\{c_{n}\}$ using Eq. (5). Alternatively, one can view $N$, $\{[f_{\mathrm{l},j},f_{\mathrm{h},j}]\}$, and $\{w_{j}\}$ as hyperparameters controlling the pulse shape.
@@
To implement fast, low-leakage single-qubit gates, we combine the FAST method with DRAG by using FAST to shape the in-phase envelope of the control pulse and applying DRAG to obtain the quadrature component. The resulting FAST DRAG pulse illustrated in Fig. 1(b) has multiple benefits compared to using DRAG with traditional pulse envelopes, such as the raised cosine. First, the FAST DRAG pulse enables a stronger and wider spectral suppression of the $ef$-transition compared to conventional approaches as shown in Fig. 1(d), which helps to mitigate leakage to the $f$-state. To achieve this, we apply the FAST method to suppress the spectrum of the in-phase component across a frequency interval $[f_{\mathrm{l},1},f_{\mathrm{h},1}]=[f_{\mathrm{l},\mathrm{ef}},f_{\mathrm{h},\mathrm{ef}}]\ni|\alpha|/(2\pi)$, and use DRAG-L with $\beta\approx 1$ to further suppress the Fourier transform $\hat{\Omega}_{\mathrm{IQ}}$ of the complex envelope that is of the product form $\hat{\Omega}_{\mathrm{IQ}}(f)=[1-2\pi\beta f/\alpha]\times\hat{\Omega}_{I}(f)$. The strong suppression of $\hat{\Omega}_{\mathrm{IQ}}(f)$ around $\alpha/(2\pi)$ then leads to a low spectral amplitude $|\hat{\Omega}_{\mathrm{RF}}(f)|$ around the $ef$-transition frequency assuming $\omega_{\mathrm{d}}=\omega_{\mathrm{q}}$. In addition to suppressing the signal around the $ef$-transition, we use a second frequency interval $[f_{\mathrm{l},2},f_{\mathrm{h},2}]$ to reduce the spectral energy above a given cutoff frequency $f_{\mathrm{c}}=f_{\mathrm{l},2}$. This favorably reduces the bandwidth of the control pulse and often indirectly lowers the peak amplitude, thus effectively regularizing the pulse shape. As a further benefit, it is relatively straightforward to calibrate FAST DRAG pulses since the parameters $A$, $\beta$, and $\{[f_{\mathrm{l},j},f_{\mathrm{h},j}]\}$ are mostly decoupled, and the locations of the suppressed frequency intervals $\{[f_{\mathrm{l},j},f_{\mathrm{h},j}]\}$ do not affect the effective drive frequency due to a symmetric suppression of the spectrum (see Fig. 1(c)).
In this work, we select the hyperparameters $\{N,f_{\mathrm{l},\mathrm{ef}},f_{\mathrm{h},\mathrm{ef}},f_{\mathrm{c}},w_{\mathrm{ef}}=w_{1}/w_{2}\}$ using physics-inspired heuristics based on the measured anharmonicity $\alpha/(2\pi)$ without optimizing the hyperparameters. In Appendix F, we provide general heuristic guidelines for the hyperparameter selection in systems with a single dominant leakage transition and discuss an extension to systems with multiple leakage transitions important for simultaneous single-qubit gates. The heuristic hyperparameter selection is possible because the leakage error is relatively robust against moderate changes in the hyperparameters as discussed in Sec. III and Appendix J. After the hyperparameter selection, the remaining single-qubit gate parameters can be calibrated similarly to conventional DRAG pulses, which renders the calibration of FAST DRAG pulses efficient and robust avoiding the challenges with closed-loop and open-loop optimal control methods mentioned in Sec. I.